# Interplay between topology and interactions in superconducting chains

### v1

@@ IV The one-dimensional spinless Falicov-Kimball model ($\Delta=0$)
We now consider the case where the conduction band has a $p$-wave superconducting pairing and interacts with the background of localized $f$-electrons. The simple, non-interacting case with $J=0$, corresponds to a Kitaev chain and an independent collection of localized electrons. The topological properties of the Kitaev chain are well known, it presents a non-trivial topological phase, for $|\mu/2t|<1$, with Majorana modes at its ends [1]. In the present case the chemical potential $\mu=0$, such that the Kitaev chain is in a topological superconducting phase [1]. We now turn on the interaction with the localized $f$-electrons and study its effect on the topological properties of the chain using simulations on finite chains.
Figure 4: Plot of the density of states in the superconducting state. a) for the uniform case for different values of the coupling parameter $J$, below, at, and above the topological transition. b) for the uniform case at the topological transition, for different pairing amplitudes $\Delta$. c) for the staggered configuration as a function of $J$ for fixed $\Delta$.
Figure 5: (Color online) Eigenvalues $\lambda/t$ for a chain of 300 sites for $t=1$, $\Delta/t=0.2$, a) in the uniform phase, for $J/t=0.1<1$, b) in the staggered phase for $J/t=0.1<\Delta/t$. In c) and d) we show the amplitude square of the wave functions of the zero modes, i.e, eigenvalues closest to zero, for the uniform and staggered phases, ( a) and b) above), respectively.
Figure 6: (Color online) Eigenvalues $\lambda/t$ for a chain of 300 sites for $t=1$, $\Delta/t=0.2$, a) in the uniform phase, for $J/t=1.3>1$, b) in the staggered phase for $J/t=1.3>\Delta/t$. Both phases are gapped and topological trivial superconductors as indicated by the absence of zero energy surface modes.
Figure 7: (Color online) Ground state energy of the $p$-wave superconductor as a function of the interaction $J$ for $\Delta/t=0.5$. Uniform case, $n_{f}=1$ (red), staggered, $n_{f}=1/2$ (blue) and 200 random configurations (green) with $<n_{f}>=1/2$. The staggered configuration has an energy lower than all the probed random configurations with $<n_{f}>=1/2$.
Figure 8: a) Eigenvalues $\lambda/t$ for $\Delta=0.2$, corresponding to 200 realizations of random configurations, with probability $p=1/2$ for a site being occupied or unoccupied by a localized electron ($<n_{f}>=1/2$) for a chain with 40 sites. a) for $J=0.1$, b) for $J=0.3$, c) for $J=1$ and d) for $J=1.2$ The zero modes persist all the way to the topological phase transition of the uniform case at $J/t=1$.
First, we discuss the results of Fig. 1 for the dispersion relations of the infinite Kitaev chain coupled to the local moments. We consider both, the uniform and staggered configurations, for some relevant set of parameters. We take, $t=1$, so that the hopping sets the energy scale for the problem, and for the pairing gap we take $\Delta/t=0.2$. In the uniform case, for $J/t=1$ there is a gap closing, and the system is a gapless superconductor. For $J\neq 1$ the superconductor is gapped, as can be seen directly from the dispersion relations and density of states in Figs. 4a. However, there is a fundamental difference between the cases $J/t<1$ and $J/t>1$, as we discuss below for finite chains.
Figure 9: a) Amplitude square of the wave functions of the lower energy modes of a chain with 180 sites for a particular realization of disorder. a) $J=0.1$. b) $J=0.2$. c) Amplitude of the wave-function of the ordered staggered configuration with $J=0.1$. Compare with a) and notice that in the random case the wave function is more localized.
@@
Fig. 4b shows the density of states for the homogeneous case in the gapless case ($J=1$), as a function of the $p$-wave superconductor parameter $\Delta$. For $\Delta$ close to zero, the FK strong correlations produce a huge peak at $\mu=0$. As the pairing interaction is turned on, a competition between the localization due to FK interactions and the formation of $p$-wave pairs occurs; the peak decreases as the superconducting pairs are formed, as indicated in the inset of the figure, and the density of states approaches that of a renormalized uncorrelated tight-binding chain. For comparison, we also show a tight-binding density of states for $J=\Delta=0.0001t$.
@@
In the staggered configuration, gapless superconductivity occurs at $J=\Delta$ and again the superconductor is gapped whenever $J\neq\Delta$ as shown in Fig. 4c. Notice in Fig. 4c that there is a finite density of states at around $\mu=0$ for $J=\Delta$, i.e., at the topological transition of the staggered configuration where the system is gapless. This is similar to the corresponding density of states of an armchair nano-ribbon, whenever $N=3M-1$, with $M$ being integer [34]. As in the uniform case there is an essential difference between the gapped phases, in the staggered phase for $J>\Delta$ and $J<\Delta$, as concerns their topological properties.
@@
In order to investigate the topological properties of the interacting system, instead of calculating a topological invariant [31, 32, 33] and using the bulk-boundary correspondence principle [34], we directly investigate the behavior of interacting finite chains using open boundary conditions. In this approach the appearance of zero energy surface states gives direct evidence of the non-trivial topological character of the system.
@@
In Fig. 5 we show the eigenvalues for a chain of 150 sites for both uniform and staggered configurations. In the case of the uniform chain, the superconductor is gapped with zero energy modes that persist, as long as $J<1$. These zero modes are surface states as can be seen directly from their wave functions shown in the same figure. At $J=1$ the gap closes and the system is a gapless superconductor. For $J>1$, the gap reopens, as shown in Fig. 1, but the zero energy surface states disappear, indicating that the superconductor is topologically trivial. Notice that in the absence of pairing, the FK chain is metallic for $J<1$, as shown in Fig. 2a. These results show that in the uniform case there is a topological phase transition, with a gap closing, at $J/t=1$.
@@
In the staggered case, the topological superconducting phase, characterized by the presence of zero energy surface states, exists for $J<\Delta$, as shown in Fig. 5b. At $J=\Delta$ the superconductor is gapless and above this critical value, the superconducting chain is gapped and topologically trivial with no zero modes. In this staggered case, as we turn on the superconducting pairing, non-trivial topological superconductivity emerges from an insulating phase (see Fig. 2d of the FK model, as soon as, the pairing amplitude becomes larger than $J$. Then, the topological phase transition occurs for $J/t=1$ in the uniform case and for $J/\Delta=1$ in the staggered case.
In Fig. 6, we show the eigenvalues of a finite chain for values of the coupling $J$ above the topological transition, both in the uniform and staggered configurations. Notice the absence of zero modes in the gap showing that these phases are topologically trivial.
For the uniform case, with a localized electron at each side of the chain there is no room for disorder. This is different in the staggered case, with $n-f=1/2$, where the localized electrons may also be randomly distributed in the chain. In Fig. 7 we plot the ground state energies of the superconductor for the uniform, staggered and random configurations as a function of the coupling $J$. Notice that for $n_{f}=1/2$ the ordered staggered configuration has always the lower energy and is the true ground state for $<n_{f}>=1/2$.
@@
In the random case, with $<n_{f}>=1/2$, a given site has a probability $p=1/2$ of being empty or occupied by a local moment. It is interesting to explore the topological character of the random chains [37, 38, 23]. Fig. 8 shows the eigenvalues for a chain with 40 sites and 200 realizations of disorder. In this case we assume that a given site has a probability $p=1/2$ of being empty or occupied by a local moment. This corresponds to taking $<n_{f}>=1/2$. We observe the presence of two zero energy surface states on average for the whole range of couplings, $J\leq 1$, as shown in Fig. 8. The wave functions of these zero modes are shown in Fig. 9 showing that these are really surface states. Compared to the ordered staggered configuration, where topological superconductivity occurs for $J_{c}=\Delta$, we conclude that disorder extends the range of topological superconductivity to much larger values of the coupling. This is a surprising effect that has been observed previously in similar systems [23].
The critical value $(J/t)_{c}=1$ that we observe numerically for the disordered system with $p=1/2$, is the same as for the homogeneous case and different from the ordered staggered configuration with the same average number of localized electrons per site, $<n_{f}>=1/2$. Notice that can write the interaction between localized and conduction electrons at a given site $i$ of the chain, as $J_{i}=Jx_{i}$ where $x_{i}=\pm 1$. Consider a random occupation of the local moments in the chain, given by, $P(x_{i})=p\delta(x_{i}-1)+(1-p)\delta(x_{i}+1)$ for a given site $i$. Since the $x_{i}$s are independent random variables the critical value $(J/t)_{c}=1$ for the topological transition is a consequence of the invariance of the Hamiltonian under the exchange $J\rightarrow-J$.
VI Conclusions and perspectives
In this work, we have studied the properties of a $p$-wave superconducting chain in the presence of an interaction between the conduction electrons and a background of localized electrons. The interaction is of the FK type and couples the charge degrees of freedom of both systems, itinerant and localized electrons. We have shown that the many-body problem can be mapped in a non-interacting one with a site dependent chemical potential, which allows it to be soluble. The solution provides a clear picture of the effects of interactions in the topological properties of the chain. We find that in the case of a chain coupled to a uniform configuration of localized electrons, i.e., with every site of the chain occupied by a localized electron, the interaction between the two systems leads eventually to the suppression of the topological phase. However, the critical value of this coupling is large, of the order of the bandwidth of the conduction electrons. In the case the localized electrons are in a staggered ordered configuration, in which every other site in the chain is occupied, the critical value of the coupling to suppress the topological phase is much reduced, being given by the pairing energy of the conduction electrons. In this case superconductivity arises from an insulating state of the FK aboce a critical value of the pairing amplitude, $\Delta_{c}=J$. This staggered configuration corresponds to $<n_{f}>=1/2$, and for most of the configurations with this concentration, the localized electrons randomly occupy the sites in the chain. We have numerically studied the case of random configurations with probability $p=1/2$ for a site in the chain being occupied or not by a localized electron, which corresponds to $<n_{f}>=1/2$. We find that randomness extends the range of stability of the topological superconducting phase with respect to the coupling to the localized electrons. While in the ordered staggered configuration, the topological superconducting phase exists for $J<\Delta$, in the presence of disorder it is stable up to $J=t$, the critical coupling of the uniform configuration. We find however that the ordered staggered configuration has always the lower energy. The exact results of the model allows to put on firm grounds the role of interaction and disorder in the topological properties of an archetypal system, namely the Kitaev chain. The model, can be extended to take into account different occupations for localized electrons with $p\neq 1/2$. Also, we can assume that the energy of the localized electrons coincide with the chemical potential, i.e., $E_{f}=\mu$, but relax the condition in the chemical potential so that $\mu\neq 0$. In this case we can study the phase diagram for different occupations of the band of conduction electrons.
@@ VI Conclusions and perspectives
An interesting challenge is to include hybridization between the localized and itinerant electrons, which gives rise to the quantum Falicov-Kimball model [28].
We could also investigate the thermoelectric properties of those systems at the topological quantum phase transition of the homogeneous ($t=\Delta$) and staggered ($J=\Delta$, cases [39]. Another possibility is to extend include spin in the model and study the spin polarization of MZMs in nanoribbons [18].

### v2

@@ V.1 Analysis of the dispersion relations and density of states
The topological properties of the Kitaev chain are well known, it presents a non-trivial topological phase, for $|\mu/2t|<1$, with Majorana modes at its ends [1]. In the present case the chemical potential $\mu=0$, such that the Kitaev chain is in a topological superconducting phase [1]. We now turn on the interaction with the localized $f$-electrons and study its effects.
@@
We start discussing the results of Fig. 1 for the dispersion relations, Eqs. 11 and 12, and the density of states obtained from these dispersions shown in Fig. 4. We consider both, the homogeneous and staggered configurations and take $\Delta/t=0.2$ in the figures. In the homogeneous case, the gap closes at $J/t=1$, as can be seen in the density of states shown in Fig. 4a. For $J\neq 1$ the superconductor is gapped, as can be obtained directly from the dispersion relations and as shown in the density of states in Fig. 4a. However, there is a fundamental difference between the cases $J/t<1$ and $J/t>1$, as we discuss below using numerical results in finite chains. Fig. 4b shows the density of states at the gap closing point of the homogeneous case for different values of the coupling $\Delta$. As the pairing interaction is turned on, a competition between the localization due to FK interactions and the formation of $p$-wave pairs occurs; the peak decreases as the superconducting pairs are formed, as indicated in the inset of the figure, and the density of states approaches that of a renormalized uncorrelated tight-binding chain. For comparison, we also show in Fig. 4b the tight-binding density of states for $J/t=\Delta/t=0.0001t$.
Figure 4: Density of states in the superconducting state. a) for the homogeneous case with $\Delta/t=0.2$ and different values of the coupling parameter $J/t$, below, at, and above the gap closing transition. b) for the homogeneous case at the gap closing transition and different pairing amplitudes $\Delta/t$. The inset shows the density of states close to $\omega=0$. c) for the staggered configuration as a function of $J/t$ and fixed $\Delta/t=0.2$.
@@
In the staggered configuration, the gap closing occurs at $J=\Delta$ and the superconductor is gapped, whenever $J\neq\Delta$, as shown in Fig. 4c for the density of states. Notice that there is a finite density of states at $\mu=0$ for $J=\Delta$, where the system is gapless. This is similar to the corresponding density of states of an armchair nano-ribbon, whenever $N=3M-1$, with $M$ being integer [34]. Like in the homogeneous case, in the staggered phase the gapped phases, for $J>\Delta$ and $J<\Delta$ are distinct with respect to their topological properties, as we discuss next.
V.2 Topological properties of finite chains
@@
In order to investigate the topological properties of the interacting system, instead of calculating a topological invariant [35, 36, 38] and using the bulk-boundary correspondence principle [34], we opt to study numerically the behavior of interacting finite chains with open boundary conditions. In this approach the appearance of zero energy edge states gives direct evidence of the non-trivial topological character of the system.
Figure 5: (Color online) Eigenvalues $\lambda/t$ for a chain of 300 sites for $t=1$, $\Delta/t=0.2$, a) in the homogeneous phase, for $J/t=0.1<1$, b) in the staggered phase for $J/t=0.1<\Delta/t$. In c) and d) we show the amplitude square of the wave functions of the corresponding zero modes for the homogeneous and staggered phases, respectively.
@@
In Fig. 5 we show the eigenvalues for a chain of 300 sites for both homogeneous and staggered configurations. In the homogeneous chain, Fig. 5a, superconductivity is gapped, with zero energy modes that persist, as long as $J<t$. These zero modes are edge states as can be seen directly from their wave functions shown in Fig. 5c. At $J/t=1$ the gap closes and the system is a gapless superconductor [39]. For $J>t$, the gap reopens as shown in Fig. 6a, but the zero energy edge states disappear, indicating that the superconductor is topologically trivial . Recall that in the absence of pairing, the homogeneous FK chain is metallic for $J/t<1$, as shown in Fig. 2a, such that topological superconductivity arises from a metallic state. Then in the homogeneous case there is a topological phase transition together with a gap closing, at $J/t=1$.
@@
In the staggered case, the topological superconducting phase, characterized by the presence of zero energy edge states, exists for $J<\Delta$, as shown in Fig. 5b and Fig. 5d. At $J=\Delta$ the superconductor is gapless and above this critical value, the superconducting chain is gapped and topologically trivial with no zero modes, as shown in Fig. 6b. In the staggered case, as we turn on the superconducting pairing, non-trivial topological superconductivity emerges from an insulating phase (see Fig. 2d) of the FK model, as soon as, the pairing amplitude becomes larger than $J$.
Figure 6: (Color online) Eigenvalues $\lambda/t$ for a chain of 300 sites for $t=1$ and fixed $\Delta/t=0.2$, a) in the homogeneous phase, for $J/t=1.3>1$, b) in the staggered phase for $J/t=1.2>\Delta/t$. Both phases are gapped and topologically trivial superconductors as indicated by the absence of zero energy edge modes.
Then, the topological phase transition occurs for $J/t=1$ in the homogeneous case and for $J/\Delta=1$ in the staggered case.
VI Random occupation of the local moments in the chain
@@
For the homogeneous case, with a localized electron at each site of the chain there is no room for disorder. This is different for the occupation of the staggered configuration, $\langle n_{f}\rangle=1/2$, where the localized electrons may also be randomly distributed in the chain. In the random case, with quenched disorder and $\langle n_{f}\rangle=1/2$, a given site has a probability $p=1/2$ of being empty or occupied by a local moment. It is interesting to explore the topological character of the random chains [23, 25, 27]. Fig. 7 shows the eigenvalues for chains with 40 sites and 200 realizations of disorder for $p=1/2$. We can observe the presence of zero energy modes for the whole range of couplings, $J\leq 1$, as shown in Fig. 7. Fig. 8 shows the wave functions of these modes for some realizations of disorder and different values of $J/t$ revealing that, in these cases, they are really edge states. In the ordered staggered configuration, topological superconductivity occurs for $J/t<\Delta/t$. In Fig. 8, where $\Delta/t=0.2$ is fixed, we notice the existence of zero energy edge states for $J/t=0.3>\Delta/t=0.2$. Then disorder extends the range of topological superconductivity to larger values of the coupling when compared to the ordered staggered configuration with the same occupation, $\langle n_{f}\rangle=1/2$. This is a surprising effect that has been observed previously in rather similar systems [23].
Figure 7: Eigenvalues $\lambda/t$ for $\Delta/t=0.2$, corresponding to 200 realizations of random configurations for a chain with 40 sites, with probability $p=1/2$ for a site being occupied by a localized electron ($\langle n_{f}\rangle=1/2$). a) for $J/t=0.1$, b) for $J/t=0.3$, c) for $J/t=1$ and d) for $J/t=1.2$ The zero modes persist all the way to the topological phase transition of the homogeneous case at $J/t=1$.
The problem of topological phases in random chains is complex and requires a closer examination. The existence of zero modes in chains with $p=1/2$ for $(J/t)\leq 1$ does not necessarily imply the topological character of the chain in this region of the phase diagram. We find that in some configurations, with zero modes their wave-functions are not localized at the edges of the chain, but in the bulk. In order to establish the topological character of the random chains a more detailed analysis is necessary. For this purpose we consider a transfer matrix approach [35, 36, 37] that has been used in the study of localization in random chains with different types of disorder [37]. This allows to define a Lyapunov exponent that can distinguish between trivial and topological phases of the superconducting chain.
Figure 8: Amplitude square of the wave functions of the lower energy modes of a chain with 180 sites for fixed $\Delta/t=0.2$, with probability $p=1/2$ for a site being occupied by a localized electron ($\langle n_{f}\rangle=1/2$) and particular realizations of disorder. a) $J/t=0.1$. b) $J/t=0.3$. c) Amplitude of the wave-function of the ordered staggered configuration with $J/t=0.1$. Compare with a) and notice that in the random case the wave function is more localized.
The Heisenberg equation of motion for the Majorana operators $\alpha_{A}$ in the Hamiltonian, Eq. 8 can be cast in the form of a recursion relation [35, 36, 38],
$\begin{pmatrix}\alpha_{An+1}\\ \alpha_{An}\end{pmatrix}=T_{i}\begin{pmatrix}\alpha_{An}\\ \alpha_{An-1}\end{pmatrix},$
with a similar equation for the $\alpha_{B}$. The relevant, zero energy, site dependent transfer matrix is given by,
$T_{i}(E=0)=\begin{pmatrix}\frac{-2J_{i}/t}{1+\Delta/t}&\frac{-(1-\Delta/t)}{1+\Delta/t}\\ 1&0\\ \end{pmatrix}.$
Notice that we can write the interaction between localized and conduction electrons at a given site $i$ of the chain, as $J_{i}=Jx_{i}$ where $x_{i}=\pm 1$ according to the probability distribution, $P(x_{i})=p\delta(x_{i}-1)+(1-p)\delta(x_{i}+1)$ for a given site.
Figure 9: (Color online) Phase diagram of chains with probability $p$ of its sites being occupied by a local moment. The lines in the figures represent the boundaries between trivial and topological phases and correspond to the points where the Lyapunov exponent $\gamma_{r}$, changes sign. The filled regions are topologically non-trivial, with $\gamma_{r}<0$. a) for $\Delta/t=0.2$. b) $\Delta/t=0.5$. In b) the horizontal orange line is the phase boundary for $\Delta/t=1$.
Figure 10: (Color online) Phase diagram showing the critical coupling as a function of the gap amplitude for a fixed probability, $p=1/2$, of occupation of a site by a local moment. The straight, dashed line is the phase boundary of the ordered staggered configuration. Disorder extends the topological region of the phase diagram.
The ratio $r_{i}=(\alpha_{Ai+1}/\alpha_{Ai})$, as can be seen from the transfer matrix, obeys the following recursion relation [37],
$r_{i+1}=-\frac{2(J/t)x_{i}}{1+(\Delta/t)}-\frac{1-(\Delta/t)}{1+(\Delta/t)}\frac{1}{r_{i}}.$ (13)
This is a random map since the $x_{i}$ take random values in different sites. We remark, the following relation involving the $r_{i}$,
$\prod_{i=1}^{N}r_{i}=\frac{\alpha_{N+1}}{\alpha_{N}}\frac{\alpha_{N}}{\alpha_{N-1}}\frac{\alpha_{N-1}}{\alpha_{N-2}}\cdots\frac{\alpha_{1}}{\alpha_{0}}=\frac{\alpha_{N+1}}{\alpha_{0}},$ (14)
where we dropped the symbol $A$ from the operators $\alpha$. The quantity above describes the behavior of the amplitude of the wave functions of the edge modes as it penetrates in the chain. It can be used to define a Lyapunov exponent,
$\gamma_{r}=\frac{1}{N}\sum_{i=1}^{N}\ln|r_{i}|.$ (15)
We have calculated numerically this Lyapunov exponent using Eq. 13 to determine the phase diagrams of random chains with different values of $\Delta/t$, $J/t$ and $p$. The topological region of the phase diagram is characterized by negative values of the Lyapunov exponents and the phase boundary is where these change sign. We carried out $15$$\times$ $10^{4}$ iterations of Eq. 13 with $r_{0}=1$ to obtain the phase diagrams shown in Figs. 9 and 10. Notice that for $\Delta/t=1$, the Lyapunov exponent can be calculated exactly. Since the product of the $x_{i}$ can only assume the values, $\prod x_{i}=\pm 1$, we have $\gamma_{r}=\ln[|J/t|]$ independently of the probability $p$. This implies that the system is topological, whenever $J/t<1$, for any p.
We point out that between the phase boundaries for $\Delta/t=1$ (yellow line) and $\Delta/t=1/2$ (blue line) in Fig. 9b, we still observe zero modes in the spectra of eigenvalues, but the system is not topological in this region. For some of these modes we find that their wave functions are localized in the bulk of the chain. Finally, Fig. 10 gives the phase diagram for a fixed value of $p=1/2$ showing the phase boundary for the disordered chain. In this figure the straight line is the critical line for the ordered staggered configuration. Notice that disorder increases the range of the topological phase in the phase diagram.
Finally, we should mention that if we use an alternative approach, and define a Lyapunov exponent from the product of the transfer matrices [38], we obtain similar results.
VII Conclusions and perspectives
In this work, we studied the properties of the one-dimensional FK model with a $p$-wave pairing of the electrons in the conduction band. We have shown that this many-body problem can be mapped in a non-interacting one, which allows it to be exactly soluble. The solution provides a clear picture of the effects of interactions and disorder in the topological properties of the superconducting chain. The conditions for solubility are quite general, and permits to treat different numbers of localized electrons in the chain and arbitrary values of the FK interaction and superconducting pairing. In this paper we focused in a half-filled conduction band but the solution can be easily extended for general occupations.
We find that for a chain with a homogeneous configuration of localized electrons, i.e., with every site of the chain occupied by a local moment, the FK interaction leads eventually to a suppression of the superconducting topological phase, at large values of $J$ ($J_{c}=t$). In this system when turning on the pairing interaction, superconductivity arises from a metallic state of the FK model. When the localized electrons assume a staggered configuration, in which every other site in the chain is occupied, the critical value of the coupling to suppress the topological phase is much reduced, being given by the pairing energy of the conduction electrons ($J_{c}=\Delta$). In this case, superconductivity arises from an insulating state of the FK model, and we can alternatively state that a minimum critical value of the pairing amplitude, $\Delta_{c}=J$, is required to induce superconductivity in the insulator.
The staggered configuration corresponds to $\langle n_{f}\rangle=1/2$. For this occupation of local moments, most of the configurations are disordered, with the localized electrons randomly occupying the sites in the chain. We have used transfer matrices and exact recursion relations to obtain the Lyapunov exponents and show that for $p=1/2$, randomness extends the range of stability of the topological superconducting phase with respect to the coupling to the localized electrons. While in the ordered staggered configuration, the topological superconducting phase exists for $J/\Delta<1$, in the presence of disorder, it is stable up to larger values of $J$, as shown in Fig. 10.
We have also studied random configurations with an arbitrary probability $p$ for a site in the chain being occupied by a localized electron. The results obtained from the Lyapunov exponents yield the phase diagrams of the model and allows to put on firm grounds the role of interaction and disorder in the topological properties of an archetypal system, namely the Kitaev chain.
@@ VII Conclusions and perspectives
Finally, an interesting challenge is to include in this problem the hybridization between the localized and itinerant electrons, which gives rise to the quantum Falicov-Kimball model [30]. We could also investigate the thermoelectric properties of these systems at the topological quantum phase transition of the homogeneous ($t=\Delta$) and staggered ($J=\Delta$), cases  [40].